# LLaRA: Supercharging Robot Learning Data for Vision-Language Policy

### v2

@@ 6.2 Real-world robot experiments
We further conduct zero-shot generalization, fine-tuning, and joint training experiments in a novel real-world environment, in which there is a robot arm with a gripper and a fixed RGB camera positioned above the arm to collect observations (see Fig. 7). The action space is the same as the one in VIMA. As all the objects are placed on a plain surface and the camera is fixed, similar to VIMA, a linear mapping between the image coordinates and the robot action space can be estimated by calibration. For the methods that rely on object detection i.e., $\mathcal{D}$-inBC, we employed one-shot detection using OWLv2 (Minderer et al., 2024).
@@
We benchmark each policy on three tasks:
@@
where the {object} is chosen from a set of 10 plastic toys and the {angle} in T2 is chosen between 0 to 180 degrees. In all tasks, all objects are placed randomly on the table before running an episode and they never appear in the training set. A success in T1 and T3 is if the object is placed more than 50% inside the bowl or on the tray. A success in T2 is if the object appears to be clockwise rotated by the angle by visual inspection. We run 20 randomly initialized episodes for each setting and report the average success rate. A visual reference for typical task start states and success ends are provided in Fig. 8.
@@
Table 3: Real-world robot experiment results. ‘Aux’ means trained on auxiliary datasets and ‘OD’ means using an object detector (OWLv2, base-16) to describe the observation before the VLM query. In Protocol, ‘ZS’ stands for zero-shot evaluation, ‘FT’ means the model is first trained on VIMA for 2 epochs and then finetuned on Add. Data for 1 epoch. ‘JT’ means the model is trained on the combination of VIMA and Add. Data for 2 epochs.
@@
FT        RT-2 Style xArm-Det 0 0 0 0
@@
JT LLaRA (Ours) inBC xArm-Action 60 75 55 63.3
@@
Finetuning on real-world data. In the finetuning setting, we first collect two in-domain real-world datasets xArm-Det and xArm-Action (see Appendix C). Then the models from Sec. 6.1 are further tuned on the new real-world datasets for 1 epoch and the evaluation results are presented in the middle of Tab. 3. In general, finetuning outperforms other settings because it allows the model to focus more on real-world data distribution, rather than be distracted by simulated VIMA data. Auxiliary data contributes to a large boost in performance during finetuning, showing that explicitly uncovering information from the existing in-domain data can benefit the model. It seems inBC pretrained with VIMA is not trained to take advantage of such new in-domain robotics data, performing worse than zero-shot. Similar to our findings in simulation experiments, LLaRA in general benefits from pretrained VLM since LLaRA is trained on a dataset that has been organized in a conversation style that is similar to what is used to pretrain the VLM. In contrast, RT-2 Style requires a large amount of data so it suffers more from limited data in our real-world experiments.
Joint training on both simulated and real-world data. In the joint training setting, we combine both VIMA data with xArm-Det or xArm-Action and tune a VLM jointly on both simulated and real-world datasets. The full results are presented at the lower part of Tab. 3. In the joint training setting, xArm-Det is equally or more beneficial than xArm-Action, highlighting the importance of detection and localization information from the real-world setting. Auxiliary data proves most beneficial when jointly trained with xArm-Det, achieving the highest performance across all joint training settings. This observation suggests that careful exploration of auxiliary dataset usage is crucial to maximize their benefits when joint training is performed.
Using image coordinates creates an intuitive and easy-to-learn link between visual observations and robot actions. In general, LLaRA demonstrates a robust capacity for generalization. Our observations confirm that it takes much less effort and data to transfer LLaRA models pretrained on simulated VIMA data to a real-world domain, while RT-2 Style suffers from the limited data.
More generalization results. We further validate the generalization ability of LLaRA on tasks with distractors. There are three test settings and none of them is included in the simulated training set or the dataset for fine-tuning: T1 w/ distractor: where we randomly put another random toy on the table as a distractor; T1 w/ complex rephrasing: where we rephrase the task description in 10 different unseen ways using GPT-4; Unseen task T4: the task is “Weight the object using the scale then place it back on the table.” which is never included in VIMA or xArm-Action. The results are presented in Tab. 4.
@@
Protocol T1 w/ distractor (%) T1 w/ complex rephrasing (%) T4 (%)
ZS 10 0 0
FT on xArm-Action 75 40 20
Our model demonstrates robust performance on generalized tasks, achieving a 75% success rate in T1 with distractors. This suggests that our VLM retains its generalization ability even after being trained on our instruction dataset. With a small amount of in-domain vision data, it can achieve strong performance on previously unseen tasks.
@@ 7 Conclusion
We present LLaRA, a framework that turns a pretrained vision language model (VLM) into a robot policy using curated instruction tuning datasets. Firstly, we create a conversation-style instruction-tuning data using existing behavior cloning data. Instruction tuning the VLM on this data confirms the effectiveness of our framework, particularly when the training data is limited, thanks to our formulation that aligns robot actions to image coordinates. Then we construct auxiliary datasets using the same robot trajectory data through self-supervision, which proved to be further beneficial to robot learning. Experiments across several synthetic environments and robot manipulation tasks in the real world validate the effectiveness of our proposed LLaRA framework.
@@ Acknowledgments
This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government foundation. [24ZB1200, Research of Human-centered autonomous intelligence system original technology]. The authors would like to thank Bohong Chen for verifying our code implementation and reproducing experiment results, as well as Xiaodong Liu and Hanyi Yu for their valuable input.
@@ A.1.2 Build $\mathcal{D}$-inBC from inBC
A.1.3 Auxiliary datasets
@@
Table 5: Comparison between our converted instruction tuning datasets inBC and $\mathcal{D}$-inBC for a VIMA-Bench sample. The task description of the episode is in blue. The description of an object in the reference image (oracle detection results) is in magenta. The action history is in orange.
@@
Table 6: Another comparison between two converted instruction tuning datasets. In this example, the reference images in the task description depict scenes with multiple objects. The task description of the episode is in blue. The description of an object in the reference image (oracle detection results) is in magenta.
@@
Table 7: Qualitative examples of the auxiliary datasets. In this table, the examples of all the datasets take the same input image as shown in the first row. The description of an object or a scene (oracle detection results) is in magenta. The action is in orange.
@@
Table 8: (Continued) Qualitative examples of the auxiliary datasets. In this table, the examples of all the datasets take the same input image as shown in the first row. The description of an object or a scene (oracle detection results) is in magenta. The examplar is in teal.
@@ B.7 Ablation on action prompt during inference
B.8 Few-shot experiments using LLM / VLM
@@ C.3 Real-world robot dataset
Tab. 13 shows examples of the real-world datasets we collected.
Table 13: Qualitative examples of our real-world datasets xArm-Det and xArm-Action. The description of an object in the image (one-shot detection results) is in magenta. The task description of the episode is in blue.
@@
C.4 Qualitative examples on unseen real-world tasks
@@ Appendix E Raw results and prompt templates
Table 14: Results on VIMA-0.8k dataset. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset.
@@
Table 16: Results on VIMA-8k dataset. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset.
@@
Table 17: Results on VIMA-8k dataset with object detector or Oracle. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset.
@@
Table 19: Results on VIMA-80k dataset with Oracle. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset.
@@
‘Could you detail the steps needed to transform the scene shown in the image into the second scene? The second scene is provided as a collection of object bounding boxes {scene}. The format for these bounding boxes is <b>(x, y), {w, h}</b>, where x and y represent the center coordinates, and w and h are the width and height. The coordinates should be normalized to a scale of 0 to 1, starting from the top left corner.’ ‘Can you describe what actions are required to rearrange the scene shown in the image to match the second scene? The second scene is given as a set of object bounding boxes {scene}. These bounding boxes follow the format <b>(x, y), {w, h}</b>, where x and y indicate the center coordinates, and w and h represent the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’ ‘Could you list the steps necessary to modify the scene shown in the image to the second scene? The second scene is described as a collection of object bounding boxes {scene}. The bounding box format is <b>(x, y), {w, h}</b>, with x and y denoting the center coordinates, and w and h representing the width and height. The coordinates are normalized to a scale of 0 to 1, starting from the top left corner.’ ‘Can you explain what needs to be done to adjust the scene shown in the image to resemble the second scene? The second scene {scene} consists of object bounding boxes provided in the format <b>(x, y), {w, h}</b>. Here, x and y represent the center coordinates, and w and h are the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’ ‘Could you outline the necessary actions to arrange the scene shown in the image into the second scene? The second scene is defined by a collection of object bounding boxes {scene}. These bounding boxes follow the format <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h are the width and height. The coordinates start from the top left corner and should be normalized to a scale of 0 to 1.’ ‘Can you specify what needs to be done to convert the scene shown in the image into the second scene? The second scene is provided as a series of object bounding boxes {scene}. The format for these bounding boxes is <b>(x, y), {w, h}</b>, with x and y representing the center coordinates, and w and h indicating the width and height. Coordinates should be normalized from the top left corner to a scale of 0 to 1.’ ‘Could you describe the steps required to change the scene shown in the image to the second scene? The second scene is depicted as a collection of object bounding boxes {scene}. The bounding box format is <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h represent the width and height. The coordinates are normalized to a scale of 0 to 1 starting from the top left corner.’
@@
‘The image shows a scene with various objects. Now you {pick and place}, what will the scene look like? List the object bounding boxes. The format for these bounding boxes is <b>(x, y), {w, h}</b>, with x and y representing the center coordinates, and w and h as the width and height. Coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’ ‘Displayed in the image is a scene containing multiple objects. Now you {pick and place}, what will the scene look like? Write the list of object bounding boxes. The bounding box format is <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h are the width and height. The coordinates start from the top left corner and are normalized to a scale of 0 to 1.’ ‘The image illustrates a scene with various objects. Now you {pick and place}, what will the scene look like? List the object bounding boxes. The bounding boxes are formatted as <b>(x, y), {w, h}</b>, where x and y indicate the center coordinates, and w and h represent the width and height. Coordinates should be normalized from the top left corner to a scale of 0 to 1.’ ‘An image depicts a scene with multiple objects. Now you {pick and place}, what will the scene look like? Write the list of object bounding boxes. The bounding box format is <b>(x, y), {w, h}</b>, where x and y represent the center coordinates, and w and h denote the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’

### v3

@@ 6.2 Real-world robot experiments
We further conduct zero-shot generalization, finetuning, and joint training experiments in a novel real-world environment, in which there is a robot arm with a gripper and a fixed RGB camera positioned above the arm to collect observations (see Fig. 7). The action space is the same as the one in VIMA. As all the objects are placed on a plain surface and the camera is fixed, similar to VIMA,
@@
a linear mapping between the image coordinates and the robot action space can be estimated by calibration. For the methods that rely on object detection i.e., $\mathcal{D}$-inBC, we employed one-shot detection using OWLv2 (Minderer et al., 2024). We benchmark each policy on three tasks:
@@
where the {object} is chosen from a set of 10 plastic toys and the {angle} in T2 is chosen from the following set: $\{30,90,120,180\}$. In all tasks, all objects are placed randomly on the table before running an episode and they never appear in the training set. A success in T1 and T3 is if the object is placed more than 50% inside the bowl or on the tray. A success in T2 is if the object appears to be clockwise rotated by the angle by visual inspection. We run 20 randomly initialized episodes for each setting and report the average success rate. A visual reference for typical task start states and success ends are provided in Fig. 8.
@@
Table 4: Real-world robot experiment results. All models are trained on VIMA-8k unless otherwise stated. ‘Aux’ means trained on auxiliary datasets and ‘OD’ means using an object detector (OWLv2-base-16) to describe the observation before the VLM query. In Protocol, ‘ZS’ stands for zero-shot evaluation, ‘FT’ means the model is first trained on VIMA-8k for 2 epochs and then finetuned on Add. Data for 1 epoch. ‘JT’ means the model is trained on the combination of VIMA-8k and Add. Data for 2 epochs.
@@
FT RT-2 Style VIMA-8k xArm-Det 0 0 0 0
RT-2 Style VIMA-660k xArm-Det 0 0 0 0
@@
inBC + Aux (D) xArm-Det 100 95 70 88.3
inBC + Aux (D) xArm-Action 70 75 80 75
$\mathcal{D}$-inBC + OD xArm-Det 0 0 0 0
@@
JT LLaRA (Ours) inBC xArm-Det 80 75 70 75
inBC xArm-Action 60 75 55 63.3
inBC + Aux (D) xArm-Det 75 85 70 76.6
inBC + Aux (D) xArm-Action 50 90 50 63.3
@@
Finetuning on real-world data. In the finetuning setting, we first collect two in-domain real-world datasets xArm-Det and xArm-Action (see Appendix C). Then the models trained on VIMA-8k introduced in Sec. 6.1 are further tuned on the new real-world datasets for 1 epoch and the evaluation results are presented in the middle of Tab. 4. In general, finetuning outperforms other settings because it allows the model to focus more on real-world data distribution, rather than be distracted by simulated VIMA data. Auxiliary data contributes to a large boost in performance during finetuning, showing that explicitly uncovering information from the existing in-domain data can benefit the model. It seems inBC is not trained to take advantage of such new in-domain robotics data, performing worse than zero-shot. Similar to our findings in simulation experiments, LLaRA in general benefits from pretrained VLM since LLaRA is trained on a dataset that has been organized in a conversation style that is similar to what is used to pretrain the VLM. In contrast, RT-2 Style requires a large amount of data so it suffers more from limited data in our real-world experiments.
Joint training on both simulated and real-world data. In the joint training setting, we combine both VIMA-8k with xArm-Det or xArm-Action and tune a VLM jointly on both simulated and real-world datasets. The full results are presented at the lower part of Tab. 4. In the joint training setting, xArm-Det is equally or more beneficial than xArm-Action, highlighting the importance of detection and localization information from the real-world setting. Auxiliary data proves most beneficial when jointly trained with xArm-Det, achieving the highest performance across all joint training settings. This observation suggests that careful exploration of auxiliary dataset usage is crucial to maximize their benefits when joint training is performed.
@@
Protocol ZS FT on xArm-Action
T1 w/ distractor (%) 10 75
T1 w/ complex rephrasing (%) 0 40
T4 (%) 0 20
More generalization results. We further validate the generalization ability of LLaRA on tasks with distractors. There are three test settings and none of them is included in the simulated training set or the dataset for finetuning: T1 w/ distractor: where we randomly put another random toy on the table as a distractor; T1 w/ complex rephrasing: where we rephrase the task description in 10 different unseen ways using GPT-4; Unseen task T4: the task is “Weight the object using the scale then place it back on the table.” which is never included in VIMA or xArm-Action. The results are presented in Tab. 6.2. Our model demonstrates robust performance on generalized tasks, achieving a 75% success rate in T1 with distractors. This suggests that our VLM retains its generalization ability even after being trained on our instruction dataset. With a small amount of in-domain vision data, it can achieve strong performance on previously unseen tasks.
Summary. Our findings highlight the strong generalization ability of LLaRA. When pretrained on simulated VIMA data, LLaRA requires significantly less effort and data to adapt to real-world tasks. By leveraging image coordinates, LLaRA establishes a clear, straightforward mapping between visual inputs and robot actions. In contrast, RT-2 Style struggles under limited data conditions, largely because of the extensive data needed to learn the semantics of newly introduced tokens.
@@ 7 Conclusion
We present LLaRA, a framework that turns a pretrained vision language model (VLM) into a robot policy using curated instruction tuning datasets. Firstly, we create a conversation-style instruction-tuning data using existing behavior cloning data and align the robot action to image pixel coordinates. Instruction tuning the VLM on such data confirms the effectiveness of our framework, particularly when the training data is limited. Then we construct auxiliary datasets using the same robot trajectory data through self-supervision, without any additional action labels. Experiments across several synthetic environments and robot manipulation tasks in the real world validate the effectiveness of our proposed LLaRA framework. We believe the new observations on the VLM to Vision-Language-Action Model (VLA) representation transfer we share with this paper are valuable not only to the robot learning community but also to the computer vision community.
@@ Acknowledgments
This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government foundation [24ZB1200, Research of Human-centered autonomous intelligence system original technology]. The authors would like to thank Bohong Chen for verifying our code implementation and reproducing experiment results, as well as Xiaodong Liu and Hanyi Yu for their valuable input.
@@ A.1.2 Build $\mathcal{D}$-inBC from inBC
A.1.3 $\mathcal{D}$-inBC (L): describing the reference images in natural language
Following the suggestion from the reviewer, we trained a variant of $\mathcal{D}$-inBC, which completely removes all the coordinate numbers in the prompt even from the reference images. Instead, we use a pretrained VLM, Qwen2-VL-7B-Instruct (Wang et al., 2024), to generate natural language descriptions of all reference images, omitting any use of image coordinates. These textual descriptions replaced the original lists of image coordinates in the input prompt for the $\mathcal{D}$-inBC setting, resulting in a variant we refer to as $\mathcal{D}$-inBC (L), where (L) stands for language. Then we train on VIMA-8k and the conversations describing the reference image. During inference, LLaRA first generates textual descriptions of all reference images using its own capabilities and subsequently leverages the generated text to perform the task.
The following prompt is used to generate image description: ‘‘Write short descriptions of each object in the image, specifying their color, texture, shape, and relative positioning.’’
Tab. 10 presents an example of generated image description of the {frame_0} in Tab. 9. We present the performance of $\mathcal{D}$-inBC (L) at Tab. 6 (VIMA) and Tab. 7 (Real-world).
Table 6: $\mathcal{D}$-inBC (L) performance when trained on VIMA-8k
Method L1 (%) L2 (%) L3 (%)
inBC + Aux (D) 59.2 58.8 52.1
$\mathcal{D}$-inBC + Aux (D) 62.3 57.7 52.1
$\mathcal{D}$-inBC (L) + Aux (D) 61.9 54.2 55.4
Table 7: $\mathcal{D}$-inBC (L) real-world experiment results.
Method Add. Data T1 (%) T2 (%) T3 (%) Avg. (%)
inBC + Aux (D) xArm-Det 100 95 70 88.3
inBC + Aux (D) xArm-Action 70 75 80 75
$\mathcal{D}$-inBC (L) + Aux (D) xArm-Det 65 90 70 75
$\mathcal{D}$-inBC (L) + Aux (D) xArm-Action 75 90 75 80
In general, $\mathcal{D}$-inBC (L) achieves comparable but slightly worse performance than $\mathcal{D}$-inBC, which uses structured text to describe the reference images.
A.1.4 Auxiliary datasets
@@
Table 8: Comparison between our converted instruction tuning datasets inBC and $\mathcal{D}$-inBC for a VIMA-Bench sample. The task description of the episode is in green. The description of an object in the reference image (oracle detection results) is in cerulean. The action is in orange and the action history is in magenta.
@@
Table 9: Another comparison between two converted instruction tuning datasets. In this example, the reference images in the task description depict reference images with multiple objects. The task description of the episode is in green. The description of an object in the reference image (oracle detection results) is in cerulean and the action is in orange.
@@
Table 10: Image description used in $\mathcal{D}$-inBC (L), an example of {frame_0} in Tab. 9.
The image displays three distinct objects, each with unique colors, shapes, textures, and spatial relationships.
1. Object on the Left:
- Color: Red
- Texture: Matte
- Shape: Triangular
- Spatial Relationship: Positioned to the far left of the image.
2. Object in the Middle:
- Color: Green and white
- Texture: Speckled
- Shape: Arrowhead
- Spatial Relationship: Situated to the right of the first object.
3. Object on the Right:
- Color: Multicolored (red, blue, green, and yellow)
- Texture: Glossy
- Shape: Trapezoidal
- Spatial Relationship: Located to the far right of the image.
These objects are arranged in a linear sequence, with each subsequent object placed to the right of the previous one, creating a horizontal alignment. The colors and textures vary, adding visual interest to the composition.
Table 11: Qualitative examples of the auxiliary datasets. In this table, the examples of all the datasets take the same input image as shown in the first row. The description of an object or a scene (oracle detection results) is in blue. The action is in orange.
@@ 3. Object on the Right:
Table 12: (Continued) Qualitative examples of the auxiliary datasets. In this table, the examples of all the datasets take the same input image as shown in the first row. The description of an object or a scene (oracle detection results) is in blue. The examplar is in purple.
@@ B.7 Ablation on action prompt during inference
B.8 Extended experiments on RT-2 Style variants
As suggested by the reviewer, we evaluate two new variants of the RT-2 Style settings, RT-2 Style (2R) and RT-2 Style + Aux (initially reported in Tab. 3). In the original RT-2 Style setting, a single token is assigned to present the rotation in the action space. However, considering there are only 256 possible values of a token, it causes an unfair comparison between our method and RT-2 Style in terms of action space of rotation and may compromise RT-2 Style’s performance. Therefore, we propose RT-2 Style (2R), which uses two tokens to represent rotation, resulting in 65,536 options to encode rotation. This new configuration introduces a significantly larger action space compared to our method. At the same time, we evaluate another variant RT-2 Style + Aux (D) by simply mixing RT-2 Style training data with auxiliary datasets used in LLaRA. Evaluation results of these models on VIMA-8k are summarized in Tab. 16.
Table 16: Performance of RT-2 Style variants on VIMA-8k
Method L1 (%) L2 (%) L3 (%)
RT-2 Style 3.8 3.1 1.7
RT-2 Style (2R) 1.9 2.3 1.7
RT-2 Style + Aux (D) 36.9 35.8 32.1
inBC 57.3 46.2 42.9
inBC + Aux (D) 59.2 58.8 52.1
As shown in Tab. 16, an additional token for rotation in RT-2 Style (2R) does not significantly change the performance of the RT-2 Style baseline, however, the introduction of auxiliary data significantly improves the performance of RT-2 Style, demonstrating the effectiveness of our method.
B.9 Robustness of the model in failure grasp scenarios
As suggested by the reviewer, to show the robustness of our method, especially the ability to recover from failure, we run additional tests in a revised VIMA environment. The revised environment has a 10% chance that ignores the action from the model to simulate the action failure, which we believe is much higher than the actual case.
Tab. 17 shows that without any changes, our existing model is robust to such failure.
Table 17: Robostness evaluation results of $\mathcal{D}$-inBC + Aux (B) + Oracle (VIMA-80k, 8 epochs)
Prob. of Failure L1 (%) L2 (%) L3 (%)
0 90.0 88.1 79.2
10 87.7 86.2 75.8
B.10 Few-shot experiments using LLM / VLM
@@ C.3 Real-world robot dataset
Table 19: Qualitative examples of our real-world datasets xArm-Det and xArm-Action. The description of an object in the image (one-shot detection results) is in blue. The task description of the episode is in green and the action is in orange..
@@
Tab. 19 shows examples of the real-world datasets we collected.
C.4 Extended real-world experiments
In addition, following the suggestion from the reviewer, we have added experiments with a new set of tasks that require more language understanding and spatial reasoning. These tasks include:
  • T5: Put all yellow objects into the bowl. There is a bowl and 3 objects on the table. Between 1 and 3 of the objects are yellow. A success state is one in which all yellow objects are in the bowl and no non-yellow objects are in the bowl.
  • T6: Transport the baked goods to the large bowl. There is a bowl and 3 objects on the table. Either 1 or 2 of the objects are baked goods. A success state is one in which all baked goods are in the large bowl and none of the non-baked goods are in the large bowl.
  • T7: Feed the pet healthy food. There is an animal on the table and 2 objects. One of the objects is a fruit or vegetable and the other is a baked good. The 2 objects are equidistant from the animal to start. A success state is one in which the healthy food is sitting next to the animal while the unhealthy food is further away from the animal.
Without a proper language understanding, these tasks cannot be easily solved. These tasks are particularly challenging as none of these tasks were seen in the training data in either sim or real. Some of the objects in the scene (e.g., baked goods and bowl) were never seen by the model either during the robot data training, and each episode takes multiple steps to finish.
We compare RT-2 Style and LLaRA (inBC + Aux (D)) both first trained on VIMA-8k and then finetuned on xArm-Det, which is the model used in Tab. 4 without any changes. That is, the new tasks and the unseen tasks for the model. For each task, we run 10 randomly initialized episodes and report the success rate. Tab. 20 shows the results.
Table 20: Real-world experiment results of more unseen tasks that require challenging semantic understanding.
Method Add. Data T5 (%) T6 (%) T7 (%)
RT-2 Style xArm-Det 0 0 0
inBC + Aux (D) xArm-Det 60 55 50
As we are able to observe, LLaRA is able to handle tasks requiring spatial language understanding tasks to a certain degree. It sometimes gets confused due to its smaller LLM size (7B), but it shows the potential to effectively leverage the knowledge of LLaVA, enabling efficient transfer to robot control with minimal data. This capability is primarily facilitated by our key contributions: the alignment between action and image coordinates, and the integration of self-supervised auxiliary data.
RT-2 Style on the other hand, simply fails to utilize the limited amount of training data provided in our setting, despite using the same model architecture.
C.5 Qualitative examples on unseen real-world tasks
@@ Appendix E Raw results and prompt templates
Table 21: Results on VIMA-0.8k dataset. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset. The ‘*’ means the reference images that appeared only in the task description are not used to generate this dataset.
@@
Table 23: Results on VIMA-8k dataset. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset. The ‘*’ means the reference images that appeared only in the task description are not used to generate this dataset.
@@
Table 24: Results on VIMA-8k dataset with object detector or Oracle. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset. The ‘*’ means the reference images that appeared only in the task description are not used to generate this dataset.
@@
Table 26: Results on VIMA-80k dataset with object detector or Oracle. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset.
@@
‘Could you detail the steps needed to transform the scene shown in the image into the second scene? The second scene is provided as a collection of object bounding boxes {scene}. The format for these bounding boxes is <b>(x, y), {w, h}</b>, where x and y represent the center coordinates, and w and h are the width and height. The coordinates should be normalized to a scale of 0 to 1, starting from the top left corner.’
‘Can you describe what actions are required to rearrange the scene shown in the image to match the second scene? The second scene is given as a set of object bounding boxes {scene}. These bounding boxes follow the format <b>(x, y), {w, h}</b>, where x and y indicate the center coordinates, and w and h represent the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’
‘Could you list the steps necessary to modify the scene shown in the image to the second scene? The second scene is described as a collection of object bounding boxes {scene}. The bounding box format is <b>(x, y), {w, h}</b>, with x and y denoting the center coordinates, and w and h representing the width and height. The coordinates are normalized to a scale of 0 to 1, starting from the top left corner.’
‘Can you explain what needs to be done to adjust the scene shown in the image to resemble the second scene? The second scene {scene} consists of object bounding boxes provided in the format <b>(x, y), {w, h}</b>. Here, x and y represent the center coordinates, and w and h are the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’
‘Could you outline the necessary actions to arrange the scene shown in the image into the second scene? The second scene is defined by a collection of object bounding boxes {scene}. These bounding boxes follow the format <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h are the width and height. The coordinates start from the top left corner and should be normalized to a scale of 0 to 1.’
‘Can you specify what needs to be done to convert the scene shown in the image into the second scene? The second scene is provided as a series of object bounding boxes {scene}. The format for these bounding boxes is <b>(x, y), {w, h}</b>, with x and y representing the center coordinates, and w and h indicating the width and height. Coordinates should be normalized from the top left corner to a scale of 0 to 1.’
‘Could you describe the steps required to change the scene shown in the image to the second scene? The second scene is depicted as a collection of object bounding boxes {scene}. The bounding box format is <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h represent the width and height. The coordinates are normalized to a scale of 0 to 1 starting from the top left corner.’
@@
‘The image shows a scene with various objects. Now you {pick and place}, what will the scene look like? List the object bounding boxes. The format for these bounding boxes is <b>(x, y), {w, h}</b>, with x and y representing the center coordinates, and w and h as the width and height. Coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’
‘Displayed in the image is a scene containing multiple objects. Now you {pick and place}, what will the scene look like? Write the list of object bounding boxes. The bounding box format is <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h are the width and height. The coordinates start from the top left corner and are normalized to a scale of 0 to 1.’
‘The image illustrates a scene with various objects. Now you {pick and place}, what will the scene look like? List the object bounding boxes. The bounding boxes are formatted as <b>(x, y), {w, h}</b>, where x and y indicate the center coordinates, and w and h represent the width and height. Coordinates should be normalized from the top left corner to a scale of 0 to 1.’
‘An image depicts a scene with multiple objects. Now you {pick and place}, what will the scene look like? Write the list of object bounding boxes. The bounding box format is <b>(x, y), {w, h}</b>, where x and y represent the center coordinates, and w and h denote the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’